The standard cost of Product B manufactured by Pharrell Company includes 3.5 units of direct materials at $5.10 per unit. During June, 27,300 units of direct materials are purchased at a cost of $4.90 per unit, and 27,300 units of direct materials are used to produce 7,700 units of Product B.

(a) Compute the total materials variance and the price and quantity variances.

Total materials variance $enter a dollar amount select an option

Materials price variance $enter a dollar amount select an option

Materials quantity variance $enter a dollar amount select an option

(b) Compute the total materials variance and the price and quantity variances, assuming the purchase price is $5.15 and the quantity purchased and used is 26,700 units.

Total materials variance $enter a dollar amount select an option

Materials price variance $enter a dollar amount select an option

Materials quantity variance $enter a dollar amount select an option
Answer #1

a) Total material variance = (7700*3.5*5.10)-(27300*4.90) = 3675 F

Material price variance = (5.10-4.90)*27300 = 5460 F

Material quantity variance = (7700*3.5-27300)*5.1 = 1785 U

b) Total material variance = (7700*3.5*5.10)-(26700*5.15) = 60 U

Material price variance = (5.10-5.15)*26700 = 1335 U

Material quantity variance = (7700*3.5-26700)*5.1 = 1275 F